OpenStudy (anonymous):

find the value of the expression 9P3

OpenStudy (raden):

use this formula : nPr = n!/(n-r)! with n! = n(n-1)(n-2)(n-3)... 3.2.1

OpenStudy (raden):

in this case, known the value of n=9 and r=3

OpenStudy (anonymous):

i have to show my work how do i do tht

OpenStudy (raden):

9P3 = 9!/(9-3)! = 9!/6! = ... simplify

OpenStudy (anonymous):

is it 504

OpenStudy (raden):

yup that's right

OpenStudy (anonymous):

thank you

OpenStudy (raden):

you're welcome

OpenStudy (anonymous):

okay simplify this expression 10C4